OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

Multiple the entire first equation by 2 and what do you get?

OpenStudy (anonymous):

You can either use elimination or subsitution. What method are u suppossed to use?

OpenStudy (anonymous):

elemination

OpenStudy (anonymous):

Alrighty so I will explain the concept

OpenStudy (anonymous):

In order to eliminate a variable they must have the same coefficient in front

OpenStudy (anonymous):

So lets eliminate the y

OpenStudy (anonymous):

Are the coeeficients in front of both y's the same?

OpenStudy (anonymous):

Are u there? lets eliminate the x since it may be easier to understand

OpenStudy (anonymous):

So do both x's have he same coefficient?

OpenStudy (anonymous):

This is a harder one since both these equations are really one

OpenStudy (anonymous):

4x-2y=12 4x-12=2y 2x-6=y Or it can be written as 2x-y =6

OpenStudy (anonymous):

So in essence they are the same equation so there are an infinitely amount of solution where both of these equations meet since they are the same line!!

OpenStudy (anonymous):

Did you get that?